# Dynamical correlations and nonequilibrium sum rules in photodoped Hubbard ladders

### v2

@@ Abstract
Using matrix product state techniques we study the nonequilibrium dynamical response of the half-filled Hubbard ladder when subject to an optical pump. Optical pumping offers a way of producing and manipulating new strongly correlated phenomena by suppressing existing magnetic correlations. The ladder allows the effects of pump directionality to be investigated, and compared to a single chain it has strong spin-charge coupling and a fully gapped excitation spectrum, promising different nonequilibrium physics. We compute time-dependent correlations, including the nonequilibrium dynamical structure factors for spin and charge. By deriving a combined spin-charge sum rule that applies both in and out-of-equilibrium, we show that spectral weight is pumped directly from the antiferromagnetic spin response into a low energy $\omega\sim 0$ charge response below the Mott gap. The transfer of weight is pump direction dependent: pumping directed along the legs disrupts magnetic correlations more than pumping in the rung direction, even if the post pump energy density is similar. The charge correlation length is dramatically enhanced by the pump, whilst the spin correlations are most strongly suppressed at nearest and next-nearest neighbour spacings. After the pump the system is in a nonthermal correlated metallic state, with gapless charge excitations and approximately equal spin and charge correlation lengths, emphasising the importance of treating these degrees of freedom on an equal footing in nonequilibrium systems.
@@ I Introduction
The dynamical response of a many-body system is a sensitive probe of the energy and momentum dependence of its collective excitations. For equilibrium systems, the evolution of the dynamical response on tuning a parameter (such as temperature, doping, or an applied field) can reveal subtle phase transitions and the behaviour of emergent collective phenomena, and act as a precise test of theoretical models [1]. In nonequilibrium condensed matter systems, the dynamical response can be interrogated in real time, thanks to the advent of ultrafast experimental probes, including time resolved resonant inelastic X-ray scattering (tr-RIXS) [2, 3, 4, 5, 6, 7, 8, 9, 10, 11]. This allows for the realisation and characterisation of nonequilibrium states in systems tuned by optical pumping (‘photodoping’). These techniques have been used to demonstrate the emergence of nonequilibrium collective phenomena when an equilibrium correlated state is suppressed by pumping [12, 13].
@@
The distribution of the dynamical response is constrained by sum rules, which relate integrals of dynamic response functions - moments of spectral weight - to static quantities. A well known case occurs in many-body spin systems, for which the integrated spin response over all energies and momenta can be related to the (constant) magnitude of the local moment. Analytical methods can be verified by checking that their integrated dynamics saturate this spin sum rule [14]. Moreover, as the total spin spectral weight is conserved, if the spin response is reduced in one region of energy-momentum space by tuning, then there is a concomitant increase in spin response elsewhere. For example, in gapped spin systems with hard core excitations, increasing temperature leads to suppression of interband spectral weight, and the emergence of a low energy intraband response due to scattering from a thermal population of spin excitations [15, 16, 17].
@@
In this work we analyse the transfer of dynamical response from the spin channel to the charge channel for simulated optical pumping of the half filled Hubbard ladder (a system of two coupled 1D Hubbard chains). The Hubbard model provides a minimal theoretical description of the interplay between spin and charge degrees of freedom in a many-body quantum system. At half filling (one electron per site) the Hubbard model is a Mott insulator (an insulator driven by interaction rather than band structure effects), with strong antiferromagnetic correlations characterising the equilibrium state.
@@
Here we emphasise that the combined total spin and charge spectral weight is a conserved (static) quantity even for the nonequilibrium Hubbard model. This combined sum rule can be used to understand that optical pumping of the ladder does not redistribute spectral weight within the spin and charge channels independently (by creating spin and charge excitations respectively). Instead there is a net transfer of spectral weight to the charge channel, in the form of a strong response at low energies associated with intraband scattering processes in the upper Hubbard band.
We show that pump anisotropy has a strong effect on the suppression of spin correlations, with pumping along the leg direction being more disruptive than pumping along the rung direction for the same fluence. We find that the suppression of the spin response is characterised by both a decrease in the antiferromagnetic correlation length and an additional large decrease in the spin correlations at short distances, relative to equilibrium. Concomitantly, as optical pumping transfers spectral weight to the charge response, the charge correlation length grows dramatically until it is similar to the spin correlation length, leaving the system in a correlated metallic state far from the Heisenberg model limit.
This paper is organised as follows. In Section II we describe the specifics of our model and pump protocol, and some aspects of the MPS methods used including the use of conserved quantities. Section III reports our results for real time dynamics during and after different simulated pumps, including the magnetisation, local doublon/holon density and the equal time spin and charge correlation functions. In Section IV we present results for the nonequilibrium spin and charge dynamical structure factors as a function of time. We summarise and discuss our findings in Section V, and provide some further technical details in the appendices.
@@ II Model and pump protocol
[…]
@@
We are interested in systems that have strong magnetic correlations in equilibrium, so we take our initial state to be the ground state of $H_{0}$ (Eq. II) at half filling, with $t_{\perp}=t_{\parallel}=t_{h}=1$ and $U=8t_{h}$. The pump is then applied, either in the leg or rung direction, and the resulting time-dependent Hamiltonian $H(t)$ causes the state to evolve in a non-trivial manner.
We employ MPS techniques to study the ensuing nonequilibrium behaviour, computing the ground state of the equilibrium system $H_{0}$ using the density matrix renormalization group (DMRG) technique, and applying time evolving block decimation (TEBD) to evolve the system under the influence of $H(t)$. We study the thermodynamic limit, i.e. infinitely long ladders, using infinite DMRG (iDMRG) and infinite TEBD (iTEBD). Our simulations are performed using ChainAMPS [29], taking into account the conservation of the total number of electrons and the total $z$-component of spin. For leg pumps we also use parity (reflection symmetry between the two legs) to improve the numerical efficiency of the MPS representation. This is not applicable for rung pumps because they break the parity symmetry. This can be seen by applying the parity operator which entails $l=1,2\to 2,1$: for a leg pump $H(t)$ remains invariant under this operation, but for a rung pump the $\perp$ hopping parameters are complex and applying the parity operator leads to
@@
so that the parameters for hopping in the rung direction are conjugated. Breaking the parity symmetry also has consequences for charge density expectation values. For leg pumps on infinite ladders the occupation number $\langle n_{i,l}\rangle=1$ by translational invariance and parity symmetry. For rung pumps translational invariance along the legs is maintained, so $\langle n_{i,l}\rangle=\langle n_{j,l}\rangle$, but the broken parity symmetry during the pump means that in general $\langle n_{i,1}\rangle\neq\langle n_{i,2}\rangle$.
Following earlier work on the Hubbard chain [18, 20], we choose a pump frequency $\Omega=6$ that is close to resonance with the Hubbard interaction scale $U=8$ (approximately the gap between the lower and upper Hubbard bands), so that significant magnetic and charge effects should be observed in the nonequilibrium state. We fix the pump width to be $\sigma_{p}=1$, centred at $t_{p}=\pi/2$, and apply the pump from $t=0$ to $t=\pi$. These values result in the pump switching on and off when its instantaneous amplitude is zero, so that the effects of an abrupt change in $H$ are diminished and we can improve numerical efficiency by using a time independent Hamiltonian for the majority of the evolution.
@@ III Dynamics in real time and space
In this section we consider real space and equal time correlations for the photodoped Hubbard ladder. We examine local observables (the energy density, magnetization, and doublon/holon density) and the spatial correlation functions for spin and charge.
@@ III.1 Local observables
In order to compare the effects of pumping along the legs versus along the rungs, we first determine the energy imparted to the system. We define $\Delta E$ as the difference between the energy after the pump ends and the energy of the initial equilibrium state (the ground state of Eq. II at half filling). Fig. 1 shows the increase in energy per site, $\Delta E/N$, for several different pumps applied to the Hubbard ladder as a function of $A_{0}^{2}$, which is proportional to the fluence of the pump. For fixed $A_{0}=1$, we find that the leg direction pump (denoted $\parallel$) of the infinite ladder imparts approximately $40\%$ more energy than the corresponding rung direction pump (denoted $\perp$). Adjusting the amplitude of the leg pump to $A_{0}=0.45$ results in a similar $\Delta E$ to that of the $A_{0}=1$ rung pump. Fig. 1 also includes exact diagonalization data for a 6 rung ladder with open boundary conditions, which shows similar trends to the infinite ladder. For leg direction pumping, the data show saturation of $\Delta E$ by $A_{0}\sim 0.8$, in qualitative agreement with the behaviour observed for the Hubbard chain [18]. For rung direction pumping this saturation must occur at higher fluences than for leg pumps, and hasn’t been achieved yet at $A_{0}=1$. The results in Fig. 1 suggest that a pump with fixed fluence can yield significantly different photoinjected energies depending on direction. In the following we consider the rung pump with amplitude $A_{0}=1$ as our reference, and compare it with leg pumps at different $A_{0}$ corresponding to either similar fluence or similar photoinjected energy.
Figure 1: Energy density of the Hubbard ladder, relative to its ground state, after pumps of the type described in Eq. 2 are applied along either the leg ()($\parallel$)( ∥ ) or rung ()perpendicular-to($\perp$)( ⟂ ) directions. Data are shown for infinitely long ($L=\infty$) and finite ($L=6$) ladders. The pumps parameters are $\Omega=6.0$ and $\sigma_{p}=1.0$. Lines are a guide to the eye.
@@
For half filling in equilibrium, $\langle d_{i,l}\rangle=\langle h_{i,l}\rangle$ because of the constraint $n_{i,l}=1$. However, when the parity symmetry between the legs is broken (as it is for a rung direction pump) this constraint is relaxed to $n_{i,1}+n_{i,2}=2$ and it is only true that $\sum_{l}\langle d_{i,l}\rangle=\sum_{l}\langle h_{i,l}\rangle$. We define $\Delta d=\sum_{l}\langle d_{i,l}(t)-d_{i,l}(0)\rangle/2$ as the difference between the instantaneous doublon density and the equilibrium value, averaged over the two legs.
@@
Figure 2: Upper panel: magnetization (left axis) and doublon number (right axis) as a function of time for the infinite Hubbard ladder when pumped along its legs ()($\parallel$)( ∥ ) or along its rungs ()perpendicular-to($\perp$)( ⟂ ) from $t=0$ to $t=\pi$. The pump parameters are $\Omega=6.0$ and $\sigma_{p}=1.0$. Lower panel: The change in doublon number relative to the initial state $\Delta d=\sum_{l}(\langle d_{i,l}(t)-d_{i,l}(0)\rangle)/2$, scaled by the pump amplitude squared $A_{0}^{2}$, showing collapse onto a single curve at short times. Dashed lines show the values of $\Delta E/(NUA_{0}^{2})$ for the various pumps.
@@
Photodoping during the pump decreases magnetisation and leads to oscillations in magnetisation and doublon number with double the pump frequency. For $U\to\infty$, the ground state of the half-filled, translationally invariant, Hubbard ladder has $\langle(S^{z}_{i,l})^{2}\rangle=1/4$ due to the presence of only one electron per site $\langle d_{i,l}\rangle=\langle h_{i,l}\rangle=0$. At finite $U$ the actual magnetization is reduced due to quantum fluctuations (virtually doubly occupied and vacant sites). For our parameters $\langle(S^{z}_{i,l})^{2}\rangle\approx 0.23$ and $\langle d_{i,l}\rangle\approx 0$, as seen at time $t=0$ in the upper panel of Fig. 2. Photodoping during the pump further suppresses the magnetization (though not necessarily monotonically) due to the enhanced density of doubly occupied and vacant sites. The suppression of magnetization (or equivalently the enhancement of doublon density) during the pump is accompanied by oscillations with a time scale $T=\pi/\Omega$, i.e. with a frequency double that of the pump. This oscillatory time dependence can be confirmed by a short time and small $A_{0}$ expansion which yields $\Delta d\sim A^{2}_{0}\sin^{2}\Omega t$. After the pump ends at $t=\pi$ the oscillations persist with damping, and the magnetisation either plateaus (e.g. for the leg pump with $A_{0}$) or appears to recover slowly (e.g. the leg pump with $A_{0}=0.35$).
@@
Comparison shows that the amplitude of oscillations is governed by $A_{0}$ (consistent with a small $t$ expansion) whilst the overall trend and post pump magnetization is related to the photoinjected energy. This can be seen in the lower panel of Fig. 2, which plots the change in doublon density $\Delta d$ scaled by the fluence, showing that $\Delta d/A_{0}^{2}$ is the same for all pumps at short times. However, at later times during the pump $\Delta d$ approaches $\Delta E/(NU)$ (indicated by dashed horizontal lines on the plot) which is the expected result for the change in doublon density when injecting energy $\Delta E/N$ per site into a model with no hopping. In the post pump period $\Delta d$ falls below $\Delta E/(NU)$. This can be interpreted as the creation of increasing numbers of approximately stationary doublons during the pump, followed by the decay of some of them and an associated increase in kinetic energy. The profile for the most energetic pump we consider, the $A_{0}=1$ leg pump, shows some decay of doublons before the pump ends at $t=\pi$. This qualitative difference to the other pumps is consistent with the saturation shown in Fig. 1.
@@
The leg direction pumps cause more disruption of magnetic correlations (and correspondingly a higher density doublons) than the rung pump, even if the energy transferred to the ladder is the same. The photoinjected energy for the rung pump is $\Delta E\approx 0.7$, close to that of the $A_{0}=0.45$ leg pump. However, the upper panel of Fig. 2 shows that the magnetization (and doublon) profile for the rung pump at later times is most similar to that of the leg pump with $A_{0}=0.4$, which has a lower photoinjected energy, $\Delta E\approx 0.6$.

### v3

@@ Abstract
Using matrix product state techniques we study the nonequilibrium dynamical response of the half-filled Hubbard ladder when subject to an optical pump. Optical pumping offers a way of producing and manipulating new strongly correlated phenomena by suppressing existing magnetic correlations. The ladder allows the effects of pump directionality to be investigated, and compared to a single chain it has strong spin-charge coupling and a fully gapped excitation spectrum, promising different nonequilibrium physics. We compute time-dependent correlations, including the nonequilibrium dynamical structure factors for spin and charge. By deriving a combined spin-charge sum rule that applies both in and out-of-equilibrium, we show that spectral weight is pumped directly from the antiferromagnetic spin response into a low energy $\omega\sim 0$ charge response below the Mott gap. The transfer of weight is pump direction dependent: pumping directed along the legs disrupts magnetic correlations more than pumping in the rung-direction, even if the post pump energy density is similar. The charge correlation length is dramatically enhanced by the pump, whilst the spin correlations are most strongly suppressed at nearest and next-nearest neighbour spacings. After the pump the system is in a nonthermal correlated metallic state, with gapless charge excitations and approximately equal spin and charge correlation lengths, emphasising the importance of treating these degrees of freedom on an equal footing in nonequilibrium systems.
@@ I Introduction
The dynamical response of a many-body system is a sensitive probe of the energy and momentum dependence of its collective excitations. For equilibrium systems, the evolution of the dynamical response on tuning a parameter (such as temperature, doping, or an applied field) can reveal subtle phase transitions and the behaviour of emergent collective phenomena, and act as a precise test of theoretical models [1]. In nonequilibrium condensed matter systems, the dynamical response can be interrogated in real time, thanks to the advent of ultrafast experimental probes, including time resolved resonant inelastic X-ray scattering (tr-RIXS) [2, 3, 4, 5, 6, 7, 8, 9, 10, 11]. This allows for the realisation and characterisation of nonequilibrium states in systems tuned by optical (including infrared) pumping, also known as ‘photodoping’. These techniques have been used to demonstrate the emergence of nonequilibrium collective phenomena when an equilibrium correlated state is suppressed by pumping [12, 13].
An area of significant interest is the application of ultrafast pump–probe techniques to Mott and charge-transfer insulators [14]. Ultrafast pump–probe work on these materials has mapped a consistent experimental storyline: femtosecond photodoping shifts spectral weight from interband Hubbard transitions into in-gap/intraband channels and can transiently yield metallic responses on sub-ps scales [15]. Seminal optical reflectivity experiments on halogen-bridged Ni chains have shown photoinduced collapse of the gap and the appearance of Drude-like weight within 130 fs [16]. Organic one-dimensional Mott salts then provided clean model systems: in ET-F2 TCNQ, pump–probe spectroscopy resolved a transient metallic state mediated by spin–charge separation, coherent holon–doublon dynamics, and pressure-tunable recombination, with later work identifying biexciton features [17, 18, 19]. Strong-field experiments have added a complementary axis: mid-IR/THz driving modulates the effective Hubbard U, and intense single-cycle THz fields can trigger an impulsive dielectric breakdown of the Mott phase and even sub-ps insulator–metal switching [20, 21, 22]. In parent cuprates, femtosecond absorption/reflectivity measurements have established photoinduced metallization and ultrafast carrier relaxation in Nd2CuO4 and La2CuO4, with $7$-fs probes directly capturing spin-sector dynamics [23, 24]. Momentum-resolved x-ray probes are now able visualize how photocarriers reshape correlations: tr-RIXS on Sr2IrO4 revealed rapid suppression of long-range magnetism with persistent short-range correlations [2, 10]. On the electronic structure side, tr-ARPES has followed gap collapse and in-gap band formation in spin-orbit Mott iridates from the first pump-probe spectra to more recent light-induced insulator–metal transitions [25, 26]. Complementary all-optical and x-ray studies of photodoped spin–orbit Mott magnets have quantified sub-ps demagnetization and recovery pathways, helping tie the charge and spin channels together [27].
In this work we simulate pump-probe studies on ladder materials (those in which pairs of 1D chains are coupled to produce a ladder like geometry). There now exists a wide body of work studying the out-of-equilibrium behavior in cuprate-based telephone number ladders, Sr14-xCaxCu24 (SCCO). In such ladder materials, ultrafast experiments have progressed from all-optical snapshots of carrier coherence to momentum-resolved views of light-stabilized phases. Femtosecond reflectivity on SCCO first showed reversible photo-conversion between insulating and metallic responses and even a hidden insulating state controlled by trains of sub-ps pulses, establishing that ladder carrier coherence can be steered on ultrafast timescales [28]. Building on this work, optical-pump/THz-probe measurements combined with femtosecond resonant soft-x-ray scattering then uncovered a symmetry-protected metastable electronic state: near-infrared pumping was shown to transfer holes from chain “reservoirs” to the ladders, partially melting equilibrium charge order, and yielding long-lived, nonthermal conductivity changes. Most recently, tr-RIXS at the oxygen K edge directly resolved a gapless, dispersive charge-order mode emerging in this light-induced state, with dispersion up to 0.8 eV from the charge-order wavevector and a slope comparable to the quasiparticle velocity. This constitutes momentum-resolved evidence that photodoping can endow ladder carriers with itinerant character while charge order becomes dynamically fluctuating [29].
@@
We theoretically determine the momentum and energy resolved nonequilibrium response of a half-filled Hubbard ladder (a system of two coupled 1D Hubbard chains) due to simulated pumps. A key aspect of our work is an investigation of how this response is constrained by sum rules. Sum rules relate integrals of dynamic response functions - moments of spectral weight - to static quantities. A well known case occurs in many-body spin systems, for which the integrated spin response over all energies and momenta can be related to the (constant) magnitude of the local moment. Analytical methods can be verified by checking that their integrated dynamics saturate this spin sum rule [30]. Moreover, as the total spin spectral weight is conserved, if the spin response is reduced in one region of energy-momentum space by tuning, then there is a concomitant increase in spin response elsewhere. For example, in gapped spin systems with hard core excitations, increasing temperature leads to suppression of interband spectral weight, and the emergence of a low energy intraband response due to scattering from a thermal population of spin excitations [31, 32, 33].
@@
We analyse the optical pumping induced transfer of dynamical response from the spin channel to the charge channel for the half filled Hubbard ladder. As the Hubbard model provides a minimal theoretical description of the interplay between spin and charge degrees of freedom in a many-body quantum system it is of significant theoretical interest in addition to its relevance to experiment. At half filling (one electron per site) the Hubbard model is a Mott insulator (an insulator driven by interaction rather than band structure effects), with strong antiferromagnetic correlations characterising the equilibrium state.
@@
Here we determine a nonequilibrium sum rule for single band Hubbard models, showing that the combined total spin and charge spectral weight is a conserved (static) quantity. This combined sum rule can be used to understand that optical pumping of the ladder does not redistribute spectral weight within the spin and charge channels independently (by creating spin and charge excitations respectively). Instead there is a net transfer of spectral weight to the charge channel, in the form of a strong response at low energies associated with intraband scattering processes in the upper Hubbard band. This nonequilibrium reduction in the total spin response is in stark contrast to spin only models, where the suppression of antiferromagnetic correlations requires an enhancement of the spin response at other energies and momenta.
We show that pump anisotropy has a strong effect on the suppression of spin correlations: for a given fluence, when the pump is along the leg-direction the larger number of nearest neighbours leads to enhanced energy transfer and therefore more disruption of spin correlations relative to pumping along the rung-direction. We find that the suppression of the spin response is characterised by both a decrease in the antiferromagnetic correlation length and an additional large decrease in the spin correlations at short distances, relative to equilibrium. Concomitantly, as optical pumping transfers spectral weight to the charge response, the charge correlation length grows dramatically until it is similar to the spin correlation length and the charge response becomes gapless, leaving the system in a correlated metallic state far from the Heisenberg model limit.
This paper is organised as follows. In Section II we describe the specifics of our model and pump protocol, and some aspects of the MPS methods used including the use of conserved quantities. Section III reports our results for real time dynamics during and after different simulated pumps, including the magnetisation, local doublon/holon density and the equal time spin and charge correlation functions. In Section IV we present the corresponding results for the nonequilibrium spin and charge dynamical structure factors as a function of time. We summarise and discuss our findings in Section V, and provide some further technical details in the appendices.
@@ II Model and pump protocol
$t_{\parallel}c^{\dagger}_{i,l,\sigma}c_{i+1,l,\sigma}+\text{h.c.}$ $\to t_{\parallel}e^{iA_{\parallel}(t)}c^{\dagger}_{i,l,\sigma}c_{i+1,l,\sigma}+\text{h.c.},$
$t_{\perp}c^{\dagger}_{i,1,\sigma}c_{i,2,\sigma}+\text{h.c.}$ $\to t_{\perp}e^{iA_{\perp}(t)}c^{\dagger}_{i,1,\sigma}c_{i,2,\sigma}+\text{h.c.}.$
@@
We are interested in systems that have strong magnetic correlations in equilibrium, so we take our initial state to be the ground state of $H_{0}$ (Eq. II) at half filling, with $t_{\perp}=t_{\parallel}=t_{h}=1$ and $U=8t_{h}$. The pump is then applied, either in the leg or rung-direction, and the resulting time-dependent Hamiltonian $H(t)$ causes the state to evolve in a non-trivial manner.
We employ MPS techniques to study the ensuing nonequilibrium behaviour, computing the ground state of the equilibrium system $H_{0}$ using the density matrix renormalization group (DMRG) technique, and applying time evolving block decimation (TEBD) to evolve the system under the influence of $H(t)$. We study the thermodynamic limit, i.e. infinitely long ladders, using infinite DMRG (iDMRG) and infinite TEBD (iTEBD). Our simulations are performed using ChainAMPS [45], a set of MPS algorithms developed to deal with systems of many coupled integrable chains [46, 47, 48, 49]. In the present case each rung is treated as a short ‘chain’ and the computational requirements of the MPS representation are reduced by taking into account the conservation of the total number of electrons and the total $z$-component of spin. For leg pumps we also use parity (reflection symmetry between the two legs) to improve the numerical efficiency of the MPS representation. This is not applicable for rung pumps because they break the parity symmetry. This can be seen by applying the parity operator which entails $l=1,2\to 2,1$: for a leg pump $H(t)$ remains invariant under this operation, but for a rung pump the $\perp$ hopping parameters are complex and applying the parity operator leads to
@@
so that the parameters for hopping in the rung-direction are conjugated. Thus, during and after a rung-direction pump, the system is no longer guaranteed to be in a state of definite parity.
Breaking the parity symmetry also has consequences for charge density expectation values. For leg pumps on infinite ladders the occupation number $\langle n_{i,l}\rangle=1$ by translational invariance and parity symmetry. For rung pumps translational invariance along the legs is maintained, so $\langle n_{i,l}\rangle=\langle n_{j,l}\rangle$, but the broken parity symmetry during the pump means that in general $\langle n_{i,1}\rangle\neq\langle n_{i,2}\rangle$. and it is only the total occupation of a rung that is conserved, $\langle n_{i,1}+n_{i,2}\rangle=2$.
Following earlier work on the Hubbard chain [34, 36], we choose a pump frequency $\Omega=6$ that is close to the Hubbard interaction scale $U=8$ (approximately the gap between the lower and upper Hubbard bands), so that significant magnetic and charge effects should be observed in the nonequilibrium state (we discuss the effect of altering $\Omega$ in Appendix B). We fix the pump width to be $\sigma_{p}=1$, centred at $t_{p}=\pi/2$, and apply the pump from $t=0$ to $t=\pi$. These values result in the pump switching on and off when its instantaneous amplitude is zero, so that the effects of an abrupt change in $H$ are diminished and we can improve numerical efficiency by using a time independent Hamiltonian for the majority of the evolution. As well as their relevance to previous theoretical works, these parameter choices are similar to those for studies of the ladder material SCCO, for which $t_{\parallel}\approx t_{\perp}$, $U=8t_{\parallel}$ and $t_{\parallel}$ is on the order of $10^{-1}\,\text{eV}$ [28, 29]). For example, if we set $t_{\parallel}\approx 0.2\,\text{eV}$ in our work then the pump we use has a length of $\approx 10\,\text{fs}$ and a wavelength of $\approx 1000\,\text{nm}$ (i.e. infrared).
@@ III Dynamics in real time and space
In this section we use finite and infinite MPS methods to compute equal time correlations during and after pumping, using the model described in Section II. We examine local observables (the energy density, magnetization, and doublon/holon density) and the spatial correlation functions for spin and charge in real space. The response in energy and momentum space, taking into account the finite lifetime of the probe pulse, will be considered in Section IV.
@@ III.1 Local observables
In order to compare the effects of pumping along the legs versus along the rungs, we first determine the energy imparted to the system. The final photoinjected energy $\Delta E$ is given by the difference between the energy after the pump ends and the energy of the initial equilibrium state (the ground state of Eq. II at half filling). Fig. 1 shows the photoinjected energy per site, $\Delta E/N$, as a function of $A_{0}^{2}$ (which is proportional to the fluence) for leg and rung-direction pumps. We include results for infinite ladders (using iTEBD) and finite ladders (using ED). For a given value of $A_{0}$, leg-direction pumping (denoted $\parallel$) leads to a larger photoinjected energy than the corresponding rung-direction pump (denoted $\perp$). For the range of $A_{0}$ shown there is a plateau starting at $A_{0}\sim 0.8$ (or $A_{0}^{2}\approx 0.64$) for leg-direction pumping, but not for rung-direction pumping.
Leg and rung-direction pumping should be expected to yield qualitative differences: the direct coupling between the pump and the motion of the electrons is only along the pump direction, and the redistribution of energy and momenta along the other direction occurs through the scattering of excitations due to interactions. The first (direct coupling) stage of this process is highly anisotropic, because the (infinite) leg direction supports continuous values of momenta $k_{x}$, but the (finite) rung direction admits only two discrete momenta $k_{y}$.
A short time expansion (see Appendix A) shows that at lowest order the photoinjected energy scales as $\Delta E\sim z_{\alpha}t_{\alpha}A_{0}^{2}\Omega^{2}$, where the index $\alpha=\parallel,\perp$ indicates the pump direction, and $z_{\parallel}=2$ and $z_{\perp}=1$ are the number of nearest neighbours in the leg and rung directions respectively. This scaling confirms that the photoinjected energy is initially be governed by the different spatial properties in the two directions. In the following we consider the rung pump with amplitude $A_{0}=1$ as our reference, and compare it with leg pumps at different $A_{0}$ corresponding to either similar fluence or similar photoinjected energy.
Figure 1: Energy density of the Hubbard ladder, relative to its ground state, after pumps of the type described in Eq. 2 are applied along either the leg $(\parallel)$ or rung $(\perp)$ directions. Data are shown for infinitely long ($L=\infty$) and finite ladders with $L=8$ and periodic boundary conditions (pbc) and $L=6$ and open boundary conditions (obc). The pump parameters are $\Omega=6.0$ and $\sigma_{p}=1.0$. Lines are a guide to the eye.
@@
For half filling in equilibrium, $\langle d_{i,l}\rangle=\langle h_{i,l}\rangle$ because of the constraint $\langle n_{i,l}\rangle=1$. However, when the parity symmetry between the legs is broken (as it is for a rung-direction pump) this constraint is relaxed to $\langle n_{i,1}+n_{i,2}\rangle=2$ and it is only true that $\sum_{l}\langle d_{i,l}\rangle=\sum_{l}\langle h_{i,l}\rangle$. We define $\Delta d=\sum_{l}\langle d_{i,l}(t)-d_{i,l}(0)\rangle/2$ as the difference between the instantaneous doublon density and the equilibrium value, averaged over the two legs.
@@
Figure 2: Upper panel: magnetization (left axis) and doublon number (right axis) as a function of time for the infinite Hubbard ladder when pumped along its legs $(\parallel)$ or along its rungs $(\perp)$ from $t=0$ to $t=\pi$. The pump parameters are $\Omega=6.0$ and $\sigma_{p}=1.0$. Lower panel: The change in doublon number relative to the initial state $\Delta d=\sum_{l}(\langle d_{i,l}(t)-d_{i,l}(0)\rangle)/2$, scaled by $z_{\alpha}A_{0}^{2}$ with $z_{\parallel}=2$ and $z_{\perp}=1$, showing collapse onto a single curve at short times. Dashed lines show the values of $\Delta E/(NUz_{\alpha}A_{0}^{2})$ for the various pumps. For convenience when comparing with Fig. 1, the squared values of $A_{0}$ are, $(0.45^{2})\approx 0.2$, $(0.4)^{2}\approx 0.16$ and $(0.35)^{2}\approx 0.12$.
@@
For $U\to\infty$, the ground state of the half-filled, translationally invariant, Hubbard ladder has $\langle(S^{z}_{i,l})^{2}\rangle=1/4$ due to the presence of only one electron per site $\langle d_{i,l}\rangle=\langle h_{i,l}\rangle=0$. At finite $U$ the actual magnetization is reduced due to quantum fluctuations (virtually doubly occupied and vacant sites). For our parameters $\langle(S^{z}_{i,l})^{2}\rangle\approx 0.23$ and $\langle d_{i,l}\rangle\approx 0.05$, as seen at time $t=0$ in the upper panel of Fig. 2. Photodoping during the pump further suppresses the magnetization (though not necessarily monotonically) due to the enhanced density of doubly occupied and vacant sites. The suppression of magnetization (or equivalently the enhancement of doublon density) during the pump is accompanied by oscillations with a time scale $T=\pi/\Omega$, i.e. with a frequency double that of the pump. The oscillatory time dependence is consistent with the short time expansion in Appendix A, which suggests that the change in magnetization should scale as $\sim z_{\alpha}t_{\alpha}A^{2}_{0}\sin^{2}\Omega t$. After the pump ends at $t=\pi$ the oscillations persist with damping due to many-body interactions, and the magnetisation either plateaus (e.g. the leg pump with $A_{0}=1.0$) or appears to recover slowly (e.g. the leg pump with $A_{0}=0.35$).
@@
The lower panel of Fig. 2 plots the change in doublon density $\Delta d$ divided by $z_{\alpha}A_{0}^{2}$, demonstrating that the results obey the short time scaling, collapsing onto a single curve for small $t$ compared to the characteristic hopping timescale $t_{h}^{-1}=1$. At later times interactions and higher order processes become important and the curves separate. Then $\Delta d$ approaches $\Delta E/(NU)$ (indicated by dashed horizontal lines on the plot) which is the expected result for the change in doublon density when injecting energy $\Delta E/N$ per site into a model with no hopping. In the post pump period $\Delta d$ falls below $\Delta E/(NU)$. This can be interpreted as the creation of increasing numbers of approximately stationary doublons during the pump, followed by the decay of some of them and an associated increase in kinetic energy. The profile for the most energetic pump we consider, the $A_{0}=1$ leg pump, shows some decay of doublons before the pump ends at $t=\pi$. This qualitative difference to the other pumps is consistent with the saturation shown in Fig. 1.
@@
The leg-direction pumps cause more disruption of magnetic correlations (and correspondingly a higher density doublons) than the rung pump, even if the energy transferred to the ladder is the same. The photoinjected energy for the rung pump is $\Delta E\approx 0.7$, close to that of the $A_{0}=0.45$ leg pump. However, the upper panel of Fig. 2 shows that the magnetization (and doublon) profile for the rung pump at later times is most similar to that of the leg pump with $A_{0}=0.4$, which has a lower photoinjected energy, $\Delta E\approx 0.6$.